find the y-intercept of 6x-y=-12
Answer by checkley77(12569)

Y=mX+b IS THE STANDARD LINE EQUATION WHERE m=SLOPE & b=Y INTERCEPT.
6x-y=-12
-Y=-6X-12
Y=6X+12 THUS THE Y INTERCEPT (b)=12